OpenStudy (anonymous):

What is the molar mass of carbon dioxide, CO2? (Round to four significant figures only). -43.99 g -44.00 g -28.00g

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

-44.00g is right?

OpenStudy (anonymous):

wait, so its 44.00? not 43.99?

OpenStudy (anonymous):

I don't know :/

OpenStudy (anonymous):

@chmvijay helppppppp

OpenStudy (anonymous):

haha hellllp! please

OpenStudy (anonymous):

yes plz its also included in my syllabus and i suck at it :(

OpenStudy (chmvijay):

molar mass of CO2 = atomic mass of C + 2 atomic mass of Oxygen =12.011 +2* 15.999 =44.009 =44.01=44

OpenStudy (anonymous):

@Ryaan, thanks tho :) & @chmvijay. thank you again

OpenStudy (chmvijay):

yw